OpenStudy (anonymous):

Marko had a piece of paper that was 3 fifths of a yard long he cut off 1 third???

OpenStudy (anonymous):

I am having trouble with the second question and the rest. P(A) is easily 0.25 but the rest I cant figure out

OpenStudy (nincompoop):

? 3/5 4/5 5/5

OpenStudy (♪chibiterasu):

@elvisg1993 Please don't post in other's questions, that's rude.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } - \frac{ 1 }{ 3 } = \frac{ 9 }{ 15 } - \frac{ 5 }{ 15 } = \frac{ 4 }{ 15 }\]